OpenStudy (anonymous):

A point charge q1= -4.00nC is at the point x= 0.60m , y= 0.80m , and a second point chargeq2 = +6.00nC is at the point x= 0.60m , y= 0. Calculate the magnitude of the net electric field at the origin due to these two point charges.

OpenStudy (anonymous):

note that \(\vec E_{total} = \vec E_{q1} + \vec E_{q2}\) where \(\vec E = \frac{Q}{4 \pi \epsilon_o r^2}(\cos \theta \vec a_x + \sin \theta \vec a_y) \) where for q1, \(r_1^2 =(|\vec r_1|)^2 =(0.6^2 +0.8^2)\), \(\theta_1 =\tan^{-1} \frac{0.8}{0.6}\) where for q2, \(r_1^2 =(|\vec r_2|)^2 =(0.6^2 +0^2)\), \(\theta_1 =\tan^{-1} \frac{0}{0.6}\) do you need further help?

OpenStudy (anonymous):

Thank you for help. I wasnt able to reply for some reason

OpenStudy (anonymous):

you're welcome :)